# VIPTR: A Vision Permutable Extractor for Fast and Efficient Scene Text Recognition

### 19 Jan 2024 v2

@@ 3.5 Subsampling
In order to save computational costs and extract significant features layer by layer, we added a down-sampling operationDu et al. (2022) on the height dimension after the first and third attention mixing stages of VIPTR. Using the features output from the previous Attention block, we first reshape it into an embedding of size $H\times W\times C_{i-1}$, representing the current height, width, and channel respectively. We then employ a $3\times 3$ convolution with a height dimension stride $2$ and a width dimension stride $1$, followed by LN, producing an embedding of size $\frac{H}{2}\times W\times C_{i}$. The Subsampling block halves the height while maintaining a constant width. It not only reduces the computational cost, but also compresses the height dimension to establish a multi-scale representation for each character without affecting the token layout in the width dimension. We also increase the channel dimension $C_{i}$ to compensate for the information loss.
@@ 3.6 Combining and Prediction
3.7 Architecture Variants
Firstly, as shown in Figure 2, in the 2-3 stages of VIPTR, we design the structures of Local and Global Attention to process character components circularly in two ways: ”series processing” and ”block-parallel processing-merging”, respectively, called VIPTRv1 and VIPTRv2. Second, we design several hyperparameters in VIPTR, including channel depth and number of heads per stage, number of mixing blocks. By varying them, VIPTR architectures with different number of parameters can be obtained. In summary, we constructed four typical architectures, namely VIPTRv1-T(Tiny), VIPTRv2-T, VIPTRv1-L(Large) and VIPTRv2-B(Base). Its detailed configuration is shown in Table 2.
@@ 3.7 Architecture Variants
Table 2: VIPTR architecture variants (w/o counting the rectification module and linear classifier).
4 Experiments
4.1 Datasets
For English STR task, our models are trained on two commonly used synthetic scene text datasets and a real scene dataset, i.e., MJSynth (MJ)Jaderberg et al. (2014, 2016) and SynthText (ST)Gupta et al. (2016) trainset (72k+55k), COCO-Text Gomez et al. (2017)(38062) and the trainsets union of ICDAR 2013 (IC13)Karatzas et al. (2013), ICDAR 2015 (IC15)Karatzas et al. (2015), IIIT5K Mishra et al. (2012) and Street View Text (SVT) Wang et al. (2011). Then the models are tested on six public benchmarks as follows: IC13 testset contains 857 selected regular testing images. SVT testset has 647 testing images cropped form Google Street View. IIIT5K-3000 is collected from the website and contains 3000 test images. IC15 is taken with Google Glasses without careful position and focusing, and testset we used has 1811 images. Street View Text Perspective (SVTP)Phan et al. (2013) is also cropped from Google Street View. There are 639 test images in this set. CUTE80 (CUTE) is proposed in Risnumawan et al. (2014) for curved text recognition. 288 testing images are cropped from full images by using annotated words.
For Chinese recognition task, we use the Chinese Scene and Document Dataset Yu et al. (2021). The former is a public dataset containing 509,164, 63,645 and 63,646 training, validation, and test images. A total of 5,880 different characters appear. The latter is also a public dataset, which containing 400,000, 50,000 and 50,000 training, validation, and test images. A total of 4,865 different characters appear. The validation set is utilized to determine the best model, which is then assessed using the test set.
@@ 4.1 Datasets
4.2 Implementation Details
VIPTR uses the rectification module Shi et al. (2018), where the image text is resized to $32\times 64$ for distortion correction. We use the AdamW optimizer with weight decay of $0.05$ for training. For English models, when CSWin is used by Local Attention, the input image size is set to $32\times 96$. Otherwise, it is set to $32\times 100$. The initial learning rate are set to $\frac{5}{10^{4}}\times\frac{batchsize}{2048}$. The cosine learning rate scheduler with $2$ epochs linear warm-up is used in all $49$ epochs. Data augmentation like stretch, perspective distortion, blur and Gaussian noise, are randomly performed during the training with probability $0.4$. The alphabet includes all case-insensitive alphanumerics. The maximum annotation length corresponding to the training sample is set to $25$. For Chinese models, the input size is set to $32\times 320$ and all images are resized or padded to this scale during training. The initial learning rate are set to $\frac{3}{10^{4}}\times\frac{batchsize}{512}$. The cosine learning rate scheduler with $5$ epochs linear warm-up is used in all $100$ epochs. Data augmentation was not used for training. Word accuracy is used as the evaluation metric. All models are trained by using 4 Tesla V100 GPUs on PyTorch.
4.3 Ablation Study
To better understand VIPTR, we perform controlled experiments on both IC13 (regular text) and IC15 (irregular text) under different settings. The English dataset used in the experiment consists of COCO-Text and part of the real scene training set, with a total of 45054 images and 500epochs of training. In addition, we also add aset of comparative experiments to verify the effectiveness of models on arbitrary lengths’ image texts. The Chinese Document Dataset is used for training and validation. Three test sets are prepared. Testset1 is the standard test set of Chinese Document Dataset, Testset2 (2367) is croped from real documents with image length in the range $[320,1500]$. Testset3 (638) gets from the same origin, whose image length is above 1500. For efficiency, all the experiments are carried out by using VIPTR-T without rectification module and data augmentation.
Methods English regular English unregular Avg Chinese Params Speed $\downarrow$
IC13 SVT IIIT5k IC15 SVTP CUTE Scene (M) (ms)
@@ 4.3 Ablation Study
ABINetFang et al. (2021) 97.4 93.5 96.2 86 89.3 89.2 92.6 - 36.7 51.3
VisionLANWang et al. (2021b) 95.7 91.7 95.8 83.7 86 88.5 91.2 - 32.8 28
VSTTang et al. (2021) 96.4 93.8 96.3 85.4 88.7 95.1 92.6 - 64 -
@@
MaskOCR*Lyu et al. (2022) 97.7 93.7 95.4 86.6 89 87.5 92.5 68.6 36 24.1
MaskOCRLyu et al. (2022) 96.8 94.7 95.3 87.1 89.3 90.6 92.7 68.8 100 -
TrOCRLi et al. (2023b) 98.3 93.2 91 84 91 89.6 90.2 - 558 20.8
AutoSTRZhang et al. (2020) 94.2 90.9 94.7 81.8 81.7 84 89.5 - 10.4 207
PREN*Yan et al. (2021) 94.7 92 92.1 79.2 83.9 81.3 88 - 29.1 40
VIPTRv1-T(Tiny) 96.8 93.5 93.3 87.2 87.1 87.4 91.4 68.7 5.89 4.2
VIPTRv2-T 95.8 92.6 92.6 85.4 86 85 90.3 65.8 5.1 3.3
@@
Table 5: Results on six English and Chinese scene benchmarks tested against existing methods. The model types from top to bottom are CTC-based, attention-based, LM-based, and others. * means lightweight version of the corresponding model. The speed is the average inference time on one NVIDIA Tesla V100 GPU with batchsize 4 (to align with Du et al. (2022)) for 3000 English image texts.
4.3.1 The Effective Permutation of Attention Blocks
@@
There are various ways to match the global and Local Attention blocks in each stage. In previous researchvDu et al. (2022), it was found that by placing the Local Attention block before the Global Attention block, it is beneficial to both feature mining of individual symbols and to guide the Global Attention block to focus on long-term dependency capture between symbol features, while effectively saving the inference cost of the model. The top half and the bottom half of Table 1 respectively show the performance of VIPTRv1-T and VIPTRv2-T composed of various Local and Global Attention and their permutations. $[L_{i}][L_{j}G_{s}][G_{t}]$ means for each stage, Local Attention blocks are carried out at first, and then Global Attention blocks. $[L_{i}][L_{j}//G_{s}][G_{t}]$ means in stage 2-3, Local and Global Attention blocks are carried in parallel. The $G_{1}$ means MHSA, the $G_{2}$ means ORSA, the $L_{1}$ means CSWin and the $L_{2}$ means MaSA. Others are defined similarly. As can be seen, almost every scheme gives certain accuracy improvement compared to CRNN. We believe that the improvements are attributed to the comprehensive perception of character components features by self-attention mechanism. The relatively large gains on irregular text further explains the self-attention mixing block is helpful to feature learning in complex scenarios. It is observed that $[L_{2}][L_{2}G_{2}][G_{1}]$ reports the best accuracy in the top half of Table 1 and $[L_{2}][L_{2}//G_{2}][G_{1}]$ achieves best in the bottom half of Table 1. The best one gives accuracy gains of 1.4% and 2.8% on IC13 and IC15 when compared with SVTR-T. It is observed that $[L_{1}][L_{1}//G_{2}][G_{1}]$ reports the fastest inferencing speed.
4.3.2 The Effectiveness of Positional Embeddings.
As seen in Table 3, different PEs behave slightly different in recognition accuracy. LePE scheme outperforms the default one (None) by 2.4% and 5.3% on the two datasets, indicating that it is effective especially for irregular text.
4.3.3 The Results for different lengths inputs testing
@@
As seen in Table 4, the performance of the three models on the three testsets is quite different. SVTR’s inference performance starts to collapse as the input length gets larger. VIPTR not only supports arbitrary length of image text input, but the inference performance will not decrease as the input length increases.
4.4 Comparison with State-of-the-Art
We compare VIPTR with previous works on six English benchmarks covering regular and irregular text and one Chinese scene dataset in Table 5. We first look into the result in English datasets. Even VIPTRv2-T, the smallest one of the VIPTR family, achieves highly competitive accuracy among the w/o LM ones. When further compared with LM-used ones and SVTR-L, VIPTRv1-L gains the best accuracy on IC15 among the six English benchmarks. VIPTRv1-L and VIPTRv2-B’s overall accuracy are on par with recent studiesFang et al. (2021); Tang et al. (2021); Du et al. (2022); Lyu et al. (2022). As a contrast, VIPTRv1-L and VIPTRv2-B enjoy its simplicity and runs faster. We then analyze the result on Chinese Scene Dataset. Du et al. (2022) gives the accuracy of 8 existing methods as shown in the table. Encouragingly, VIPTR performs considerably well. Compared with SVTR-L, the best one among the listed ones, accuracy gains 0.5%. Other VIPTR methods also have advanced Chinese recognition accuracy. The above results are explained as VIPTRs can better perceive multi-grained character component features. They also achieved breakthroughs in characterizing Chinese words with rich stroke patterns.
@@ 5 Conclusion
We have presented VIPTR, a customized vision permutable extractor for fast and efficient STR. It extracts multi-grained character features that describe both stroke-like local patterns and inter-component dependence of varying distance at multiple height scales. Therefore, enjoying the merits of accuracy, efficiency and cross-lingual versatility. VIPTR with different capacities are also devised to meet diverse application needs. Experiments on both English and Chinese benchmarks basically verify the proposed VIPTR. Highly competitive or even better accuracy is observed compared to state-of-the-art methods, while running faster. We hope to combine VIPTR with an efficient text decoder to further improve the performance of STR.

### 24 Jan 2024 v3

@@ 3.5 Subsampling
To save computational costs and extract significant features layer by layer, we added a down-sampling operation Du et al. (2022) on the height dimension after the first and third attention mixing stages of VIPTR. Using the output of the features from the previous Attention block, we first reshape it into an embedding of size $H\times W\times C_{i-1}$, representing the current height, width, and channel respectively. We then employ a $3\times 3$ convolution with a height dimension stride $2$ and a width dimension stride $1$, followed by LN, producing an embedding of size $\frac{H}{2}\times W\times C_{i}$. The Subsampling block halves the height while maintaining a constant width. It not only reduces the computational cost, but also compresses the height dimension to establish a multi-scale representation for each character without affecting the token layout in the width dimension. We also increase the channel dimension $C_{i}$ to compensate for the information loss.
@@ 3.6 Combining and Prediction
3.7 Vision Permutation and Architecture Variants
Firstly, as shown in Figure 2, in the 2-3 stages of VIPTR, we design the structures of Local and Global Attention to process character components circularly in two ways: ”series processing” and ”block-parallel processing-merging”, respectively, called VIPTRv1 and VIPTRv2. Second, we design several hyperparameters in VIPTR, including channel depth and number of heads per stage, and number of mixing blocks. By varying them, VIPTR architectures with different numbers of parameters can be obtained. In summary, we constructed four typical architectures, namely VIPTRv1-T (Tiny), VIPTRv2-T, VIPTRv1-L (Large), and VIPTRv2-B (Base). Its detailed configuration is shown in Table 2.
@@ 3.7 Vision Permutation and Architecture Variants
Table 2: VIPTR architecture variants (w/o counting the rectification module and CTC decoder).
4 Experiments and Discussion
Datasets
Our English STR models are trained on two synthetic datasets, MJSynth (MJ) Jaderberg et al. (2014, 2016) and SynthText (ST) Gupta et al. (2016) trainsets (72k+55k), along with real-world data from COCO-Text Gomez et al. (2017). Testing is conducted on six public benchmarks: ICDAR 2013 (IC13) Karatzas et al. (2013) testset with 857 regular images, Street View Text (SVT) Wang et al. (2011) testset with 647 images from Google Street View, IIIT5K-3000 Mishra et al. (2012) sourced from websites, ICDAR 2015 (IC15) Karatzas et al. (2015) comprising 1811 images captured via Google Glasses, Street View Text Perspective (SVTP) Phan et al. (2013) with 639 images also from Google Street View, and CUTE80 (CUTE), which focuses on curved text recognition, with 288 images.
For the Chinese recognition task, we employ the Chinese Scene and Document Dataset Yu et al. (2021). The former includes a scene text dataset with 509,164 training, 63,645 validation, and 63,646 test images, featuring 5,880 unique characters. The document text dataset comprises 400,000 training, 50,000 validation, and 50,000 test images, with 4,865 distinct characters.
Implementations
The recommended training and testing configurations for VIPTR are listed in Table 3. When Training English models, data augmentation like stretch, perspective distortion, blur, and Gaussian noise, are randomly performed with probability $0.4$. Then, It was not used when training the Chinese model. All models are trained by PyTorch on four Tesla V100 GPUs.
Setup English model Chinese model
Rectification module Shi et al. (2018)
Distortion correction $32\times 64$
Optimizer AdamW
Weight decay 0.05
Lr scheduler Cosine annealing
Input size $32\times 96$ $32\times 320$
Batch size 1024 512
Initial learning rate 5e-4×$\times\frac{batchsize}{2048}$ 3e-4×$\times\frac{batchsize}{512}$
Training epochs 49 100
Warm-up epochs 2 5
Evaluation metric Word accuracy
Table 3: Training configurations for VIPTR.
Comparison with State-of-the-Art
As indicated in Table 6, we compare VIPTR with prior studies on six English and one Chinese scene benchmarks. Even the fastest variant, VIPTRv2-T, achieves highly competitive accuracy, particularly among models without Language Models (LM). In English datasets, VIPTRv1-L outperforms LM-used models and SVTR-L, even achieving the best accuracy on IC15. Both VIPTRv1-L and VIPTRv2-B demonstrate overall accuracy comparable to recent studies Fang et al. (2021); Tang et al. (2021); Du et al. (2022); Lyu et al. (2022) while maintaining simplicity and faster execution.
Turning to the Chinese Scene Dataset, the accuracy of 8 existing methods is provided by Du et al. (2022). Encouragingly, VIPTR performs notably well, with a 0.5% accuracy gain compared to SVTR-L, the best-performing model among those listed. Other VIPTR variants also exhibit advanced Chinese recognition accuracy. These results are attributed to VIPTR’s superior ability to perceive multi-grained character component features and its breakthroughs in characterizing Chinese words with rich stroke patterns.
4.1 Further Analysis
Ablation Experiments
To gain a deeper insight into VIPTR, we conducted controlled experiments on both IC13 (regular) and IC15 (irregular) under various configurations. The English dataset used in the experiment consists of COCO-Text and part of the real scene training set, with a total of 45054 images over 500 epochs of training. In addition, we conduct a series of comparative experiments to verify the effectiveness of our models on image texts of varying lengths. The Chinese Document Dataset is used for training and validation. Three test sets are employed: Testset1, the standard set from the Chinese Document Dataset; Testset2, consisting of 2,367 images cropped from real documents with lengths between 320 and 1500 pixels; and Testset3, featuring 638 images, also derived from real documents but with lengths exceeding 1500 pixels. For efficiency, all the experiments are carried out by using VIPTR-T without the rectification module and data augmentation.
@@ 4.1 Further Analysis
Methods English Regular English Unregular Avg. Chinese Scene Params (M) Speed $\downarrow$ (ms)
IC13 SVT IIIT5k IC15 SVTP CUTE
@@
ABINet Fang et al. (2021) 97.4 93.5 96.2 86.0 89.3 89.2 92.6 - 36.7 51.3
VisionLAN Wang et al. (2021b) 95.7 91.7 95.8 83.7 86.0 88.5 91.2 - 32.8 28.0
VST Tang et al. (2021) 96.4 93.8 96.3 85.4 88.7 95.1 92.6 - 64.0 -
@@
MaskOCR* Lyu et al. (2022) 97.7 93.7 95.4 86.6 89.0 87.5 92.5 68.6 36.0 24.1
MaskOCR Lyu et al. (2022) 96.8 94.7 95.3 87.1 89.3 90.6 92.7 68.8 100.0 -
TrOCR Li et al. (2023b) 98.3 93.2 91.0 84.0 91.0 89.6 90.2 - 558.0 20.8
AutoSTR Zhang et al. (2020) 94.2 90.9 94.7 81.8 81.7 84.0 89.5 - 10.4 207.0
PREN* Yan et al. (2021) 94.7 92.0 92.1 79.2 83.9 81.3 88.0 - 29.1 40.0
VIPTRv1-T (Tiny) 96.8 93.5 93.3 87.2 87.1 87.4 91.4 68.7 5.9 4.2
VIPTRv2-T 95.8 92.6 92.6 85.4 86.0 85.0 90.3 65.8 5.1 3.3
@@
Table 6: Results on six English and Chinese scene benchmarks tested against existing methods. The model types from top to bottom are CTC-based, attention-based, LM-based, and others. * means the lightweight version of the corresponding model. The speed is the average inference time on one Tesla V100 GPU with batch size 4 (to align with Du et al. (2022)) for 3000 English image texts.
Effective Permutation of Attention Blocks
@@
There are various ways to match the global and Local Attention blocks in each stage. In previous researchv Du et al. (2022), it was found that by placing the Local Attention block before the Global Attention block, it is beneficial to both feature mining of individual symbols and to guide the Global Attention block to focus on long-term dependency capture between symbol features, while effectively saving the inference cost of the model. The top half and the bottom half of Table 1 respectively show the performance of VIPTRv1-T and VIPTRv2-T composed of various Local and Global Attention and their permutations. $[L_{i}][L_{j}G_{s}][G_{t}]$ means for each stage, Local Attention blocks are carried out at first, and then Global Attention blocks. $[L_{i}][L_{j}//G_{s}][G_{t}]$ means in the stage 2-3, Local and Global Attention blocks are carried in parallel. The $G_{1}$ means MHSA, the $G_{2}$ means ORSA, the $L_{1}$ means CSWin and the $L_{2}$ means MaSA. Others are defined similarly.
It can be observed that almost every scheme gives a certain accuracy improvement compared to CRNN. We believe that the improvements are attributed to the comprehensive perception of character components features by the self-attention mechanism. The relatively large gains on irregular text further explain the self-attention mixing block is helpful to feature learning in complex scenarios. It is observed that $[L_{2}][L_{2}G_{2}][G_{1}]$ reports the best accuracy in the top half of Table 1 and $[L_{2}][L_{2}//G_{2}][G_{1}]$ achieves best in the bottom half of Table 1. The best one gives accuracy gains of 1.4% and 2.8% on IC13 and IC15 when compared with SVTR-T. It is observed that $[L_{1}][L_{1}//G_{2}][G_{1}]$ reports the fastest inferencing speed.
4.1.1 Effectiveness of Positional Embeddings.
As indicated in Table 4, various PE schemes exhibit slight variations in recognition accuracy. The LePE scheme surpasses the default one (None) by 2.4% and 5.3% on the two datasets, demonstrating its effectiveness, particularly in the case of irregular text.
4.1.2 Results for different lengths inputs testing
@@
In Table 5, the performance of the three models on the three test sets is quite different. SVTR’s inference performance starts to collapse as the input length gets larger. VIPTR not only supports arbitrary lengths of image text input, but the inference performance will not decrease as the input length increases.
@@ 5 Conclusion
In this study, we propose VIPTR, an innovative vision permutable extractor for fast and efficient Scene Text Recognition (STR). VIPTR excels at extracting character features and capturing contextual dependencies between character components, all while being unrestricted by input size during inference. These grants VIPTR notable accuracy, efficiency, and cross-lingual adaptability. We have also developed different versions of VIPTR to cater to a range of application requirements. Experiments on both English and Chinese benchmarks generally validate the effectiveness of our models. It achieves highly competitive or superior accuracy compared to state-of-the-art methods while maintaining faster processing speeds. Our future goal is to integrate VIPTR with an efficient text decoder to further elevate STR performance.